# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v3

@@ Conclusion:
Let $P$ be the distribution of a cryptographic scheme’s output under distribution $D_{1}$, and $Q$ be the distribution under $D_{2}$. For any event $E$ (e.g., adversary succeeds):
$P(E)\leq(Q(E)+\epsilon)^{(\alpha-1)/\alpha}\cdot R_{\alpha}^{\epsilon}(P\|Q)+\epsilon$
@@
$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}(\mathcal{A})\leq\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}(\mathcal{A})+\epsilon}\cdot R_{2}^{\epsilon}+q_{S}\cdot\epsilon$
where $q_{S}$ is the number of signing queries, and $R_{2}^{\epsilon},\epsilon$ are computed below. Since $R_{2}^{\epsilon}\leq 1.05$ for $|S|\leq 33$, this gives $\mathsf{Adv}_{\mathsf{IH}}\leq 1.05\sqrt{\mathsf{Adv}_{\mathsf{U}}}+\mathsf{negl}$.
Proof Q.12.
@@
In $\Pi_{\mathsf{IH}}$ (real): $\mathbf{y}=\sum_{i\in S}\mathbf{y}_{i}$ where each $\mathbf{y}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$.
@@
For $n=|S|$ uniforms on range $N=2\lfloor\gamma_{1}/|S|\rfloor$ with shift $c=\|(c\mathbf{s}_{1})_{j}\|_{\infty}\leq\beta$:
@@
Concrete values:
$|S|$ $R_{2}^{\epsilon}-1$ Security loss
4 $8.9\times 10^{-6}$ Negligible
9 $2.3\times 10^{-4}$ Negligible
17 $2.9\times 10^{-3}$ $0.3\%$
25 $\times 10^{-2} <1.4<1.4\%$
33 $\times 10^{-2} <4.1<4.1\%$
@@
Conclusion: Applying Lemma Q.10 with $\alpha=2$ over $q_{S}=2^{64}$ signing queries with $|S|\leq 33$:
$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}$ $\leq\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+\epsilon}\cdot R_{2}^{\epsilon}+q_{S}\cdot\epsilon$
$\leq 1.05\cdot\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+10^{-30}}+2^{64}\cdot 10^{-30}$
$\leq 1.05\cdot\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}}+2^{-36}$
Security interpretation: The bound involves $\sqrt{\mathsf{Adv}_{\mathsf{U}}}$ due to the generic Rényi probability transfer. For ML-DSA-65 with $\mathsf{Adv}_{\mathsf{U}}\leq 2^{-192}$ (NIST Level 3), we have:
$\mathsf{Adv}_{\mathsf{IH}}\leq 1.05\cdot 2^{-96}+2^{-36}<2^{-35}$
This provides approximately 35-bit security margin above any polynomial-time attacker. For practical deployments where $|S|\leq 17$, the tighter bound $R_{2}^{\epsilon}\leq 1.003$ gives security margin $2^{-36}$.
Remark: For Fiat-Shamir with Aborts signatures specifically, tighter reductions exist [Raccoon2024] that avoid the square root loss by exploiting rejection sampling structure. Our conservative bound suffices for the claimed security level.
@@
$\lambda_{i}=(-1)^{i-1}\cdot\binom{T}{i}\cdot\frac{T!}{i\cdot(T-i)!}=(-1)^{i-1}\cdot\binom{T}{i}$
@@
Proof Q.15.
Standard combinatorial identity. The approximation follows from Stirling’s formula.
@@
For ML-DSA-65 with $q=8380417\approx 2^{23}$:
  • For $T\leq 22$: $\max_{i}|\lambda_{i}|<q$, so Lagrange coefficients “fit” in $\mathbb{Z}_{q}$
  • For $T\geq 23$: $\max_{i}|\lambda_{i}|\geq q$, so coefficients wrap around mod $q$
@@
Lemma Q.20 (Challenge Invertibility in $R_{q}$).
Let $q=8380417$ and $R_{q}=\mathbb{Z}_{q}[X]/(X^{256}+1)$. For a uniformly random ML-DSA challenge $c\in\mathcal{C}$ (sparse polynomial with $\tau=49$ coefficients in $\{-1,+1\}$):
$\Pr[c\text{ is invertible in }R_{q}]\geq 1-256/q>1-2^{-15}$
Proof Q.21.
@@
By a union bound over 256 evaluation points, each with failure probability $\approx 1/q$:
$\Pr[c\text{ not invertible}]\leq 256/q<2^{-15}$
Monte Carlo verification with $10^{4}$ samples confirms 100% of random challenges are invertible.

### v4

@@ 5 Output
Let $P,Q$ be distributions such that $R_{2}^{\epsilon}(P\|Q)<\infty$ with one-sided smoothing ($\epsilon$-close variant $P^{\prime}$ of $P$). For any event $E$:
$P(E)\leq\sqrt{R_{2}^{\epsilon}(P\|Q)\cdot Q(E)}+\epsilon$
where $\epsilon$ is the total variation distance between $P$ and its smoothed variant $P^{\prime}$. This follows from $P(E)\leq P^{\prime}(E)+\epsilon\leq\sqrt{R_{2}(P^{\prime}\|Q)\cdot Q(E)}+\epsilon$ by Cauchy-Schwarz.
@@
$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}(\mathcal{A})}+q_{S}\cdot\epsilon$
where $q_{S}$ is the number of signing queries, $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the full-vector Rényi divergence, and the per-coordinate $R_{2}^{\epsilon,\mathsf{coord}},\epsilon$ are computed below.
Proof Q.14.
@@
In $\Pi_{\mathsf{IH}}$ (real): $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ where each $\hat{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$ is party $i$’s constant-term contribution in the nonce DKG.
@@
For $n=|S|$ uniforms on range $N=2\lfloor\gamma_{1}/|S|\rfloor$ with shift $c=\|(c\mathbf{s}_{1})_{j}\|_{\infty}\leq\beta$, specializing to $\alpha=2$ (where the factor $\alpha(\alpha-1)/2=1$):
@@
Concrete values (per-coordinate): Since nonce coordinates are sampled independently, the full-vector Rényi divergence is $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ where $n\ell=1280$ for ML-DSA-65:
$|S|$ $R_{2}^{\epsilon,\mathsf{coord}}-1$ $R_{2}^{\mathsf{vec}}=(R_{2}^{\mathsf{coord}})^{1280}$ $\sqrt{R_{2}^{\mathsf{vec}}}$ $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$
4 $8.9\times 10^{-6}$ $\approx 1.011$ $\approx 1.006$ $\approx 2^{-96}$
9 $2.3\times 10^{-4}$ $\approx 1.34$ $\approx 1.16$ $\approx 2^{-96}$
17 $2.9\times 10^{-3}$ $\approx 2^{5.4}$ $\approx 2^{2.7}$ $\approx 2^{-93.3}$
25 $1.4\times 10^{-2}$ $\approx 2^{25}$ $\approx 2^{12.5}$ $\approx 2^{-83.5}$
33 $4.1\times 10^{-2}$ $\approx 2^{75}$ $\approx 2^{37.5}$ $\approx 2^{-58.5}$
@@
Conclusion: Applying Lemma Q.12 with the full-vector Rényi divergence $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$, and taking a union bound over $q_{S}$ queries for the smoothing cost:
$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}}+q_{S}\cdot\epsilon$
Security interpretation: For ML-DSA-65 with $\mathsf{Adv}_{\mathsf{U}}\leq 2^{-192}$ (NIST Level 3), $\epsilon<10^{-30}\approx 2^{-100}$, and $q_{S}\leq 2^{30}$:
$|S|$ $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$ Total bound
$\leq 17$ $\approx 2^{-93.3}$ $\approx 2^{-70}$ ($q_{S}\cdot\epsilon$ dominant)
$\leq 25$ $\approx 2^{-83.5}$ $\approx 2^{-70}$ ($q_{S}\cdot\epsilon$ dominant)
$\leq 33$ $\approx 2^{-58.5}$ $\approx 2^{-58.5}$ ($\sqrt{R_{2}^{\mathsf{vec}}}$ dominant)
For $|S|\leq 25$, the $q_{S}\cdot\epsilon$ term dominates and the Rényi composition cost is modest (<13<13 bits). For $|S|\geq 25$, the full-vector $\sqrt{R_{2}^{\mathsf{vec}}}$ term becomes significant due to the $(1+\delta)^{1280}$ growth.
Remark: The per-coordinate composition $(R_{2}^{\mathsf{coord}})^{n\ell}$ is a consequence of applying the Rényi game hop at the nonce sampling stage, where coordinates are independent. Tighter reductions for Fiat-Shamir with Aborts [Raccoon2024] can exploit the rejection sampling structure (which acts on the full vector $\mathbf{z}$) to potentially avoid this exponential composition. Developing such tight multi-query bounds is an interesting open problem.
@@
$\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}=(-1)^{i-1}\cdot\binom{T}{i}$
@@
Proof Q.17.
The numerator of $\lambda_{i}$ is $\prod_{j=1,j\neq i}^{T}j=T!/i$ and the denominator is $\prod_{j=1,j\neq i}^{T}(j-i)=(-1)^{i-1}(i-1)!(T-i)!$, giving $\lambda_{i}=(-1)^{i-1}T!/(i!(T-i)!)=(-1)^{i-1}\binom{T}{i}$. The approximation follows from Stirling’s formula.
@@
For ML-DSA-65 with $q=8380417$, the maximum Lagrange coefficient magnitude $\binom{T}{\lfloor T/2\rfloor}$ satisfies:
  • For $T\leq 25$: $\max_{i}|\lambda_{i}|\leq\binom{25}{13}=5{,}200{,}300<q$, so Lagrange coefficients “fit” in $\mathbb{Z}_{q}$
  • For $T\geq 26$: $\max_{i}|\lambda_{i}|\geq\binom{26}{13}=10{,}400{,}600>q$, so coefficients wrap around mod $q$
@@
Lemma Q.22 (Challenge Invertibility in $R_{q}$ (Heuristic)).
Let $q=8380417$ and $R_{q}=\mathbb{Z}_{q}[X]/(X^{256}+1)$. For a uniformly random ML-DSA challenge $c\in\mathcal{C}$ (sparse polynomial with $\tau=49$ coefficients in $\{-1,+1\}$), $c$ is invertible in $R_{q}$ with overwhelming probability.
Proof Q.23 (Justification).
@@
Heuristic bound. Since $c$ is sparse rather than uniform over $R_{q}$, the evaluation $c(\omega^{2i+1})$ is a sum of 49 terms in $\mathbb{Z}_{q}$ (each $\pm\omega^{(2i+1)j}$), which is not rigorously uniform. However, treating each evaluation as approximately uniform yields the heuristic bound $\Pr[c\text{ not invertible}]\lessapprox 256/q<2^{-15}$ via union bound.
Monte Carlo verification with $10^{4}$ independently sampled challenges confirms 100% invertibility.
Protocol-level mitigation. To avoid relying on this heuristic, the protocol includes an explicit check: if $c$ is not invertible in $R_{q}$, the signing session aborts and retries. Since non-invertibility occurs with negligible probability (none observed in $10^{4}$ trials), this adds no practical overhead.